# CVC: A Large-Scale Chinese Value Rule Corpus for Value Alignment of Large Language Models

### v4

@@ Results
As shown in Table 2, across all themes, each of the seven mainstream LLMs selected CVC-generated options in over 70.5% of cases, demonstrating broad recognition and the strong applicability of CVC rules. Notably, the "Drugs" theme exhibits the highest model agreement (up to 99.6%), indicating a high degree of alignment between model choices and CVC’s value representations. Overall, both theme-level and model-level results show a consistent preference for CVC rules: the themes of Drugs, Suicide, Firearms, and Prejudice achieve preference rates of 99.6%, 94.7%, 86.1%, and 80.6%, respectively.
Table 2: Model consistency with CVC across six moral themes
@@
Suicide $0.96_{\pm 0.006}$ $0.99_{\pm 0.004}$ $0.89_{\pm 0.011}$ $0.96_{\pm 0.005}$ $0.88_{\pm 0.010}$ $0.96_{\pm 0.002}$ $0.99_{\pm 0.004}$ 0.947
Average 0.875 0.99 0.755 0.745 0.705 0.810 0.820
Chinese LLMs (e.g., DeepSeek-V3, Doubao-1.5-256k) consistently show higher alignment with CVC than Western models (e.g., GPT-4o, Claude-3-Sonnet), suggesting that CVC better reflects locally grounded values and provides more culturally appropriate alignment guidance. This distinction is particularly evident in the "Politics" theme, where the highest preference among domestic models reaches 86%, compared to only 47% for the best-performing Western model—highlighting underlying differences in political value orientation.
Human evaluation results in Table 3 further substantiate these findings: five independent annotators exhibited over 87.5% agreement with CVC across all themes, reinforcing that CVC more accurately captures value norms within the Chinese sociocultural context.
Table 3: Human consistency with CVC across six moral themes
@@
Suicide $1.00$ $0.87$ $1.00$ $0.97$ 1.01.01.01.00 $0.968$
Average $0.960$ $0.915$982.982.982.982 $0.913$ $0.875$
4.3 Task III: CVC-Driven Moral Dilemma Generation and Evaluation
In ethics and moral psychology, moral dilemmas entail decisions between competing obligations, each involving the compromise of certain valueslemmon1962moral . The classic trolley problem exemplifies this tension between deontological and consequentialist principlesthomson1984trolley . Yet, most existing dilemma datasets are manually curated, which inherently constrains their scalability, thematic diversity, and representational depthhendrycks2020aligning ; lourie2021scruples . Crafting such scenarios is cognitively intensive and often fails to encompass the complexity of real-world value conflicts. To address these limitations, we present a rule-based generation framework built upon CVC, enabling the large-scale creation of culturally grounded, value-conflict-rich dilemmas. This approach offers a more systematic and context-sensitive means of evaluating LLMs’ value alignment and preference patterns.
@@ 4.3 Task III: CVC-Driven Moral Dilemma Generation and Evaluation
To construct such rule pairs, we adopt a two-step process: (1) We use roberta-large-mnliliu2019roberta with Sentence Transformers to calculate semantic similarity, retaining pairs with scores above 0.5 to ensure topical relevance; (2) we apply all-mpnet-base-v2song2020mpnet to assess contradiction probability, keeping only pairs above 0.8 to ensure decision-relevant conflict. This process yields 80,901 rule pairs capable of inducing value-based dilemmas. Representative examples are shown in Table 13, Appendix D.
@@
The second core of automatic moral dilemma generation is prompt design. The goal is for the LLM to generate evaluation scenarios based on the given rule pairs and simultaneously generate corresponding option sets. The generated options must meet three categories: (1) Options that align with rule 1, (2) Options that align with rule 2, and (3) Distractor options. To avoid model bias toward certain options, the content of options A, B, and C is randomized, meaning that A does not always correspond to rule 1, B to rule 2, etc. After comparing the generation results of GPT-4o, Qwen2.5-72B, and DeepSeek-V3, we finally selected Qwen2.5-72B for generating evaluation scenarios. For details on prompt design and model selection comparison, please refer to Appendix D.1. In the prompt, the LLM was asked to generate five diverse scenarios using a set of conflicting rule pairs. This process resulted in 404,505 moral dilemmas. A sample of these dilemmas is presented in Figure 15 in Appendix D.1.
@@ Results
To evaluate the effectiveness of CVC-generated moral dilemmas, we randomly sampled 10,998 instances from a pool of 404,505 and tested 17 LLMs varying in origin, version, and scale (see Appendix F). The resulting similarity matrix (Figure 4(a)) reveals clear behavioral patterns: Chinese models (e.g., DeepSeek, Doubao, Qwen) show high internal consistency (typically >0.80), reflecting shared cultural orientations, while non-Chinese models form looser clusters, with some (e.g., GPT-4o, Gemini, Llama-3) showing partial alignment. Cross-cultural comparisons (e.g., Claude-3-Sonnet, Codestral vs. Chinese models) yield lower similarity (0.69–0.76), highlighting the influence of cultural and institutional factors. Notably, 7,191 dilemmas elicited divergent responses across models, demonstrating CVC’s ability to generate diverse and challenging scenarios. Moreover, the distribution of option preferences (Figure 4(b)) shows balanced choices around 50%, indicating well-constructed value trade-offs without systematic bias. Further analyses are provided in Appendix D.2 and D.3.
@@
5 Conclusion
In this work, we present the Chinese Values Corpus (CVC), a large-scale dataset containing over 250,000 human-annotated rules across national, societal, and individual dimensions, covering 12 core and 50 derived values. The proposed CVC can effectively guide the generation of value assessment scenarios, demonstrating significant advantages in value boundaries and content diversity. Furthermore, CVC accurately reflects the universal values embedded in existing LLMs, with high-quality human annotation ensuring consistent adaptation and alignment with Chinese values. Compared to other benchmarks, it provides more localized and precise standards for value alignment. Beyond single-value assessment, CVC facilitates the generation of moral dilemmas, providing a novel avenue for exploring value preferences in complex scenarios. Its potential for cross-cultural value alignment underscores its broader applicability in global ethical assessments.
Acknowledgments and Disclosure of Funding
Use unnumbered first level headings for the acknowledgments. All acknowledgments go at the end of the paper before the list of references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found at: https://neurips.cc/Conferences/2025/PaperInformation/FundingDisclosure.
Do not include this section in the anonymized submission, only in the final paper. You can use the ack environment provided in the style file to automatically hide this section in the anonymized submission.
Appendix A Chinese Values Corpus
@@ A.1 Conceptual Interpretation
Table 6: Distribution of sources for basic scenario data
Source Data Volume
Zhihu-KOL 4,847
People’s Daily 8,840
Flames 671
Chinese Moral Sentence Dataset 28,536
Encyclopedia QA (JSON version) 11,622
Web Crawling 178,056
@@ A.5 Rule Attributes
Building on the comparison results in Appendix A.4, we continue to use Qwen2.5-72B for the structured attribute classification of the CVC. The complete prompt is shown in Figure 8.
@@ A.6 Quality Control
We recruited 45 annotators through online postings. Among them, 40 annotators with backgrounds in artificial intelligence or philosophy participated in the annotation of CVC, while the remaining 5 were involved in the annotation tasks for Experiment 4.1. All annotators signed informed consent forms prior to their participation. The average annotation time was approximately 3 hours for the 40 primary annotators, and 1.5 hours for the 5 annotators in Experiment 4.1. Compensation was provided at 1.5 times the local minimum hourly wage. Before annotation, the 40 primary annotators received comprehensive training, which included an introduction to the annotation task and illustrative examples. The detailed training content is outlined below.
@@
The attribute statistics of the CVC retained after annotation by both human annotators and LLMs are shown in the Figure 10.
Figure 10: Distribution of data across the three value levels in the CVC.
Appendix B CVC-Guided Scenario Generation
@@
Appendix C Chinese Value Alignment of CVC vs. Existing Benchmarks
In this section, we present value rules extracted from three different rule corpus. For each theme, the number of rules selected from CVC is equal to the total number of rules selected from SC101 and MIC combined. All pairs of value rules are ensured to be relevant to the corresponding theme. The detailed rule selection information is shown in Table 11. The evaluation scenarios for the six themes are distributed as shown in Table 12.To minimize potential value priming, annotators received only a single instruction: “Given a scenario and its associated set of options, select the most appropriate course of action from your perspective.” This minimal guidance was intended to ensure that responses reflected the annotators’ own judgments rather than externally imposed value preferences.
Table 11: Rule pairs extracted from CVC and SC101/MIC
Theme Rule from CVC Rule from SC101/MIC
@@
Table 12: Number of evaluation scenarios per theme
Theme Number of Scenarios
Surrogacy 35
Prejudice 5
Politics 50
Firearms 35
Drugs 15
Suicide 30
Appendix D CVC-Driven Moral Dilemma Generation and Evaluation
@@ D.1 Dilemma Scenarios Creation
Table 13: Similarity and collision values between pairs of moral rules
Rule 1 Rule 2 Similarity Collision
Respect everyone. You should respect dogs. 0.5601 0.9914
@@ D.3 Case Study
Dilemma 1: Survival Instinct vs Moral Duty
@@
Dilemma 2: Child Autonomy vs Parental Authority
@@
Dilemma 3: National Responsibility vs Global Humanitarianism
@@
Appendix E Limitations
This section briefly outlines the current limitations of our study. First, regarding the use of the CVC, we have primarily applied it to guide the generation of evaluation scenarios, particularly those involving challenging ethical themes. In the future, we plan to further explore its potential as an evaluator and use it to support value alignment in LLMs. Second, in the moral dilemma analysis presented in Section 4.3, we only conducted case-based analysis due to the large data scale. A more comprehensive, global analysis of model behavior is planned for future work, along with a systematic categorization of moral dilemma types.
@@ Appendix F Model Cards
DeepSeek DeepSeek-R1[50] DeepSeek-R1
DeepSeek DeepSeek-V3[40] DeepSeek-V3
@@
Alibaba Qwen2.5-7B[36] Qwen2.5-7B-Instruct
Alibaba Qwen2.5-32B[36] Qwen2.5-32B-Instruct
Alibaba Qwen2.5-72B[36] Qwen2.5-72B-Instruct
@@
Zhipu AI GLM-4-32B[51] GLM-4-32B-0414
OpenAI GPT-4o[39] gpt-4o
OpenAI O1[52] o1
OpenAI GPT-3.5-Turbo[53] gpt-3.5-turbo-1106
@@
Google Gemini-1.5-Pro[45] gemini-1.5-pro
Meta Llama-3-70B[44] aihubmix-Llama-3-1-70B-Instruct
Meta Llama-3-405B[44] aihubmix-Llama-3-1-405B-Instruct

### v5

@@ Results
As shown in Table 4, across all themes, each of the seven mainstream LLMs selected C-VARC generated options in over 69.6% of cases (weighted average), demonstrating broad recognition and the strong applicability of C-VARC rules. Notably, the "Drugs" theme exhibits the highest model agreement (up to 99.6%), indicating a high degree of alignment between model choices and C-VARC’s value representations. Overall, both theme-level and model-level results show a consistent preference for C-VARC rules: the themes of Drugs, Suicide and Prejudice achieve preference rates of 99.6%, 94.7% and 80.6%, respectively.
Table 4: Model consistency with C-VARC across six moral themes
@@
Suicide $6_{\pm 0.006} 0.99±0$$9_{\pm 0.004}$ $0.89_{\pm 0.011}$ $0.96_{\pm 0.005}$ $0.88_{\pm 0.010}$ $6_{\pm 0.002} 0.99±0$$9_{\pm 0.004}$ 0.947
Average 0.864 0.974 0.742 0.738 0.696 0.801 0.815
Chinese LLMs (e.g., DeepSeek-V3, Doubao-1.5-256k) consistently show higher alignment with C-VARC than Western models (e.g., GPT-4o, Claude-3-Sonnet), suggesting that C-VARC better reflects locally grounded values and provides more culturally appropriate alignment guidance. This distinction is particularly evident in the "Politics" theme, where the highest preference among domestic models reaches 86%, compared to only 47% for the best-performing Western model—highlighting underlying differences in political value orientation.
Human evaluation results in Table 5 further substantiate these findings: five independent annotators exhibited over 87.5% agreement with C-VARC across all themes, reinforcing that C-VARC more accurately captures value norms within the Chinese sociocultural context.
Table 5: Human consistency with C-VARC across six moral themes
@@
Suicide $1.00$ $0.87$ $1.00$ $0.97$ 1.01.00 $0.968$
Average $0.960$ $0.915$ $0.982$ $0.913$ $0.875$ -
4.3 Task III: C-VARC Driven Moral Dilemma Generation and Evaluation
In ethics and moral psychology, moral dilemmas entail decisions between competing obligations, each involving the compromise of certain valueslemmon1962moral. The classic trolley problem exemplifies this tension between deontological and consequentialist principlesthomson1984trolley. Yet, most existing dilemma datasets are manually curated, which inherently constrains their scalability, thematic diversity, and representational depthhendrycks2020aligning; lourie2021scruples. Crafting such scenarios is cognitively intensive and often fails to encompass the complexity of real-world value conflicts. To address these limitations, we present a rule-based generation framework built upon C-VARC, enabling the large-scale creation of culturally grounded, value-conflict-rich dilemmas. This approach offers a more systematic and context-sensitive means of evaluating LLMs’ value alignment and preference patterns.
@@ 4.3 Task III: C-VARC Driven Moral Dilemma Generation and Evaluation
To construct such rule pairs, we adopt a two-step process: (1) We use roberta-large-mnliliu2019roberta with Sentence Transformers to calculate semantic similarity, retaining pairs with scores above 0.5 to ensure topical relevance; (2) we apply all-mpnet-base-v2song2020mpnet to assess contradiction probability, keeping only pairs above 0.8 to ensure decision-relevant conflict. This process yields 80,901 rule pairs capable of inducing value-based dilemmas. Representative examples are shown in Table 6.
Table 6: Similarity and collision values between pairs of moral rules
Rule 1 Rule 2 Similarity Collision
Respect everyone. You should respect dogs. 0.5601 0.9914
@@
The second core of automatic moral dilemma generation is prompt design. The goal is for the LLM to generate evaluation scenarios based on the given rule pairs and simultaneously generate corresponding option sets. The generated options must meet three categories: (1) Options that align with rule 1, (2) Options that align with rule 2, and (3) Distractor options. To avoid model bias toward certain options, the content of options A, B, and C is randomized, meaning that A does not always correspond to rule 1, B to rule 2, etc. After comparing the generation results of GPT-4o, Qwen2.5-72B, and DeepSeek-V3, we finally selected Qwen2.5-72B for generating evaluation scenarios. For details on prompt design and model selection comparison, please refer to Appendix D.1. In the prompt, the LLM was asked to generate five diverse scenarios using a set of conflicting rule pairs. This process resulted in 404,505 moral dilemmas. A sample of these dilemmas is presented in Figure 6.
Figure 6: Examples of moral dilemmas.
@@ Results
To evaluate the effectiveness of C-VARC generated moral dilemmas, we randomly sampled 10,998 instances from a pool of 404,505 and tested 17 LLMs varying in origin, version, and scale (see Appendix F). The resulting similarity matrix (Figure 7(a)) reveals clear behavioral patterns: Chinese models (e.g., DeepSeek, Doubao, Qwen) show high internal consistency (typically >0.80), reflecting shared cultural orientations, while non-Chinese models form looser clusters, with some (e.g., GPT-4o, Gemini, Llama-3) showing partial alignment. Cross-cultural comparisons (e.g., Claude-3-Sonnet, Codestral vs. Chinese models) yield lower similarity (0.69–0.76), highlighting the influence of cultural and institutional factors. Notably, 7,191 dilemmas elicited divergent responses across models, demonstrating C-VARC’s ability to generate diverse and challenging scenarios. Moreover, the distribution of option preferences (Figure 7(b)) shows balanced choices around 50%, indicating well-constructed value trade-offs without systematic bias. Further analyses are provided in Appendix D.2 and D.3.
Summary
The Chinese Value Rule Corpus (C-VARC) provides a large-scale, high-quality dataset containing over 250,000 human-annotated rules across national, societal, and individual dimensions, encompassing 12 core and 50 derived values. The dataset enables consistent and diverse generation of value assessment scenarios, demonstrating clear advantages in value coverage and content diversity. Empirical evaluations further confirm that C-VARC can effectively represent the value orientations embedded in existing large language models, offering a reliable and localized reference for assessing value alignment in Chinese-language contexts.
@@
Data Availability
The Chinese Value Rule Corpus (C-VARC) is publicly available at https://huggingface.co/datasets/Beijing-AISI/C-VARC. The dataset contains over 250,000 value rules categorized across three dimensions, twelve core values, and fifty derived values. All data are provided in structured JSON Lines format to facilitate reproducibility and integration into downstream applications, including scenario generation and value alignment evaluation. A subset of 10,000 value rules has also been translated into English for cross-linguistic research purposes. Comprehensive documentation, data schema, and usage examples are included in the repository. The dataset is fully anonymized and released for academic and non-commercial use.
Code Availability
The code used for data generation and analysis is available at https://github.com/Beijing-AISI/C-VARC. In this repository, we provide a detailed explanation of the data pipeline, including the steps for data collection, processing, and rule generation. The repository also contains instructions for replicating the data generation process and using the dataset.
Funding
This work was supported by the Beijing Major Science and Technology Project under Contract (Grant No. Z241100001324005) and the Beijing Natural Science Foundation (Grant No. 4252052).
Author Contributions
Co-first authors P.W., G.S., D.Z., and Y.W. contributed equally to the conception, supervision, and execution of this work, including the study design, experiments, and manuscript preparation. Y.D. provided constructive suggestions for the experiment “Chinese Value Alignment of C-VARC vs. Existing Benchmarks.” Y.S. contributed to data processing during the experimental phase. E.L. participated in project discussions and provided valuable feedback. F.Z. and Y.Z. were responsible for manuscript revision and overall quality control. All authors reviewed and approved the final manuscript.
Competing interests
The authors declare no competing interests.
Ethics Declarations
This study involves crowdsourced data annotation, and no sensitive personal information was collected during the data annotation process. Although an Institutional Review Board (IRB) approval was not sought due to the absence of sensitive data collection or potential risks to participants, an internal ethical review process was conducted. All annotators signed informed consent forms acknowledging their understanding of the task’s nature and potential risks. Participants were also made aware of their voluntary participation and were free to withdraw from the study at any time without penalty. The informed consent forms are available on the project’s GitHub repository. A completed Human Data Checklist is provided as a supplementary file in this submission.
Appendix A Chinese Value Rule Corpus
@@ A.5 Rule Attributes
Building on the comparison results in Appendix A.4, we continue to use Qwen2.5-72B for the structured attribute classification of the C-VARC. The complete prompt is shown in Figure 11.
@@ A.6 Quality Control
We recruited 45 annotators through online postings. Among them, 40 annotators with backgrounds in artificial intelligence or philosophy participated in the annotation of C-VARC, while the remaining 5 were involved in the annotation tasks for Experiment 4.1. All annotators signed informed consent forms prior to their participation. The average annotation time was approximately 3 hours for the 40 primary annotators, and 1.5 hours for the 5 annotators in Experiment 4.1. Compensation was provided at 1.5 times the local minimum hourly wage. Before annotation, the 40 primary annotators received comprehensive training, which included an introduction to the annotation task and illustrative examples. The detailed training content is outlined below.
@@
The attribute statistics of the C-VARC retained after annotation by both human annotators and LLMs are shown in the Figure 4.
Appendix B C-VARC Guided Scenario Generation
@@
Appendix C Chinese Value Alignment of C-VARC vs. Existing Benchmarks
In this section, we present value rules extracted from three different rule corpus. For each theme, the number of rules selected from C-VARC is equal to the total number of rules selected from SC101 and MIC combined. All pairs of value rules are ensured to be relevant to the corresponding theme. The detailed rule selection information is shown in Table 13. The evaluation scenarios for the six themes are distributed as shown in Table 3.To minimize potential value priming, annotators received only a single instruction: “Given a scenario and its associated set of options, select the most appropriate course of action from your perspective.” This minimal guidance was intended to ensure that responses reflected the annotators’ own judgments rather than externally imposed value preferences.
Table 13: Rule pairs extracted from C-VARC and SC101/MIC
Theme Rule from C-VARC Rule from SC101/MIC
@@
Appendix D C-VARC Driven Moral Dilemma Generation and Evaluation
@@ D.3 Case Study
Dilemma 1: Survival Instinct vs. Moral Duty
@@
Dilemma 2: Child Autonomy vs. Parental Authority
@@
Dilemma 3: National Responsibility vs. Global Humanitarianism
@@
Appendix E Future Directions
This study represents an initial effort to construct a large-scale, structured corpus of Chinese value rules and to demonstrate its applicability in guiding value-centric evaluation and task generation. Building on these foundations, several promising avenues for future research can be identified.
First, beyond its current use as a generator of culturally grounded evaluation scenarios, the corpus holds significant potential as an independent evaluator of model alignment. Future work will explore methodologies for leveraging the rule set to automatically assess LLM outputs, thereby extending its utility from task construction to alignment verification.
Second, in the analysis of moral dilemmas, this work primarily adopted case-based examinations to showcase the discriminative capacity of the framework. Future research will aim to complement this with large-scale, systematic analyses, encompassing: (1) cross-model consistency assessments across broader linguistic and cultural contexts, (2) statistical profiling of option selection patterns to detect latent biases, and (3) visualization of value preference distributions at multiple hierarchical levels. Such a holistic approach will enable deeper insights into the mechanisms of model alignment and divergence.
Third, recognizing the dynamic nature of social values, we envision establishing mechanisms for the corpus to evolve over time. While the Socialist Core Values provide a stable normative foundation, long-term alignment requires adaptability to emerging shifts in societal priorities. Future iterations will incorporate (1) periodic corpus reviews aligned with updated policy and institutional guidelines, (2) expanded annotation pipelines that integrate bottom-up inputs to capture evolving perspectives, and (3) automated monitoring using LLMs to detect early signals of changing moral discourse. These measures will support continuous versioning and ensure the corpus remains both temporally robust and culturally representative.
Together, these future directions highlight the dual role of the corpus: as a benchmark for assessing alignment and as a paradigm for generating ethically rich evaluation tasks. By systematically extending its analytical depth and adaptability, we aim to contribute to the development of culturally grounded, globally comparable frameworks for value alignment in large language models.
@@ Appendix F Model Cards
DeepSeek DeepSeek-R1[guo2025deepseek] DeepSeek-R1
DeepSeek DeepSeek-V3[liu2024deepseek] DeepSeek-V3
@@
Alibaba Qwen2.5-7B[yang2024qwen2] Qwen2.5-7B-Instruct
Alibaba Qwen2.5-32B[yang2024qwen2] Qwen2.5-32B-Instruct
Alibaba Qwen2.5-72B[yang2024qwen2] Qwen2.5-72B-Instruct
@@
Zhipu AI GLM-4-32B[glm2024chatglm] GLM-4-32B-0414
OpenAI GPT-4o[hurst2024gpt] gpt-4o
OpenAI O1[jaech2024openai] o1
OpenAI GPT-3.5-Turbo[ye2023comprehensive] gpt-3.5-turbo-1106
@@
Google Gemini-1.5-Pro[team2024gemini] gemini-1.5-pro
Meta Llama-3-70B[grattafiori2024llama] aihubmix-Llama-3-1-70B-Instruct
Meta Llama-3-405B[grattafiori2024llama] aihubmix-Llama-3-1-405B-Instruct